Given the following demand function for good X;

  Qdx=50-0.4Px+0.5Py-0.25B

Where Px is price of good X;

      Py is price of good Y;

   B is cosumer’s income.

Suppose Px=20; Py=30; and B=52

Find the value of:

Price elasticity of demand 

Cross elasticity of demand

Income elasticity of demand

What is the relationship between good X and good Y?

What would you advise the producer of good Y when the price of good X goes up? 


1
Expert's answer
2021-01-14T11:20:08-0500


price elasticity = ("\Delta"Q/"\Delta"P)* (P/Q)

Qdx = 44

"\therefore" price elasticity = -0.4 * (20/44) = -0.18


cross elasticity = ("\Delta"Qx/"\Delta"Py) * (Py/Qx)

=(0.5) * (30/44) = 0.34


income elasticity = ("\Delta"Q/"\Delta"Y) * (Y/Q)

=--0.4 * (52/44) = -0.47


Relationship between X and Y

Goods x and y are substitutes

The producer of good y should maintain the price of the good y when the price for good x goes up so as to increase sales